# Learning to crawl: benefits and limits of centralized vs distributed control

### v2

@@ II.2.2 Centralized control
We introduce centralization by defining an agent as an assembly of several suckers and springs which we call Control Center (CC), as illustrated by panel c of fig. 1. The number of CCs controlling the crawler depends on how many suckers are included in each CC and can range from several CCs down to a single CC, which corresponds to the highest degree of centralization. Centralized architectures control several suckers and springs simultaneously, hence have access to a richer representation of the dynamical state of the crawler, and can realize many more policies than those realizable in the distributed setting. The state of the CC is given by the compression/elongation state of each spring located between the first and last sucker within the CC; while the action of the CC consists of specifying the binary adhesion/no-adhesion instruction for each of the suckers within the CC. Since the CC has access to all combinations of compression/elongation and adhesion/no-adhesion states and actions, the state-action space scales exponentially as $|\mathcal{S}|=2^{n_{s}-1}$ and $|\mathcal{A}|=2^{n_{s}}$, with $n_{s}$ the number of suckers controlled by the given CC (for $n_{s}$ suckers under a CC there are $n_{s}-1$ connected springs). Note that to facilitate comparisons and have the same dimensionality of the Q-matrix for each CC we made the arbitrary choice to ignore the spring between two adjacent CCs.
II.2.3 The hive assumption
@@
Whenever multiple agents control crawling, a simplifying assumption may be adopted imposing that all agents learn the same policy. This is the so called hive assumption or Population-Based Training [17, 18, 19], where a population of agents is used to find a shared solution. In all our multiagent learning schemes –including distributed and centralized with multiple CCs– we compare standard learning to hive learning, as illustrated in fig. 1. In practice, each agent writes and reads the same Q-matrix. The distributed hive Q-matrix becomes a 8X2 tensor, where in addition to the 4 states associated to the internal sucker there are 2 extra entries for the tail, (null,elongated) and (null, compressed), and for the head, (elongated,null) and (compressed,null) [^4] Note that in this case the problem is effectively hive only for the internal suckers, whilst tail and head remain genuinely independent (since their states are unique) as in standard multiagent RL.. For the hive CCs there is no need to modify the Q-matrix between hive and standard learning, since all CCs have Q-matrices that are identical in shape, as discussed above. Clearly, the hive assumption speeds up training as it allows to parallelize the learning process. However, it also affects significantly the learned policies, as all agents are forced to learn the same policy, despite being placed at different locations relative to the propagating wave. In other words, the hive update imposes a consensus mechanism and all agent must “agree” on the same policy. In fact the removing the hive policy we find that different agents will learn distinct policies, and the hive update reduces the number of policies that are learned, as discussed in the Results section.
II.3 Learning protocol
Figure 2: Illustration of the training protocol on a 12 suckers crawler for various learning architectures. Note that when there are several agents the same procedure is applied to each agent (one Q matrix per agent). In the “exploration” phase, where several (sub-optimal) policies are stored, each color change represents a different policy (see also fig. 5). The specific value of $\alpha$ (learning rate) and $\epsilon$ used for exploring sub-optimal policies are case-specific and their order of magnitude is shown in table 1.
@@ II.3 Learning protocol
RL theory and algorithms are extensively covered in specialized books and articles (e.g. [16]). As briefly summarized before, in tabular Q-learning during training we iteratively update the Q-matrix using eq. 6. To keep a certain degree of exploration during training we adopt the standard procedure of following an $\epsilon$-greedy policy, that is, with a finite probability $\epsilon$ a random action is chosen instead of the one prescribed by the current policy (i.e. the action maximizing the current estimate of the value). Duriring training, both the exploration $\epsilon$ and learning rate $\alpha$ in eq. 6 are scheduled to decrease linearly, to ensure convergence to a final estimate of Q. Note that convergence to the optimal Q matrix is only guaranteed asymptotically, hence different trainings will lead to different sub-optimal policies, which we explore systematically in the following. In this work the scheduling process starts at a value of $\epsilon_{\max}=0.9$ and $\alpha_{\max}=0.1$ and is scheduled down to $\epsilon_{\min}=0.001$ and $\alpha_{\min}=0.001$. Inspired by standard episodic RL tasks, we call with the term episode any scheduling event of $\alpha$ and $\epsilon$. For episodic tasks, an episode is usually defined by a “termination” after which the agent is reset to another initial state and moves on to the next episode. However, the task explored in this work, which consists in achieving fast and unidirectional motion, is not episodic but continuing, since there is no termination. Therefore, here an episode is composed of an arbitrary number of integration steps with the same value of $\alpha$ and $\epsilon$. The discount rate is fixed to $\gamma=0.99$ to maximize long term reward $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}$. For continuing tasks, another possible formulation of the reward is that of average reward [16, 20], which does not require discount and where the reward becomes also an estimate. Preliminary results however do not indicate any visible benefits of this formulation in our problem. Note that although the equations of motion are deterministic, our learning problem is stochastic. There are two sources of stochasticity: i. the coarse grained state space and ii. the definition of agents. Both i. and ii. imply that the reward signal is a stochastic variable. Indeed, the speed of the crawler depends on the full dynamical state of the system, defined by the (continuous) speed, position and adhesion of all suckers and on the relaxed length of all springs. However, as per i., our crawler does not observe the full dynamical state of the system, but rather its coarse projection onto discrete binary states of compression vs elongation of all the springs. Thus given the actions of all suckers, the center of mass is not fully determined yielding a stochastic reward $R(t)$. Moreover, in all multiagent architectures each agent is unaware of other agents’ actions, as per ii., hence once again the actions of any individual agent do not individually determine the speed of the center of mass (i.e. the reward). This latter source of noise in the reward is bypassed by the fully centralized architecture with a single control center (1CC).
@@
With this in mind, to monitor progress during learning we considered the average value $G$, with $G=1/|\mathcal{S}|\sum_{s}\max_{a}\left(Q(s,a)\right)$, and the average taken over the states. As illustrated by fig. 2, we set an automatic protocol consisting in repeating the learning procedure until a plateau in $G$ is detected: when the plateau is not encountered after 1500 episodes, we repeat the process changing the scheduling by increasing the number of steps per episode. When several Q-matrices are learned simultaneously (for instance, in standard distributed control) the convergence is monitored separately for each agent. Due to the stochastic nature of the problem, we noticed that repeating the learning procedure yields different policies. Without attempting a systematic analysis of the ensemble of sub-optimal policies, we design a protocol to account for the stochasticity of this learning process. To efficiently store several sub-optimal policies, without repeating the entire training procedure, we introduced the following strategy: when convergence is detected, the learning parameters are raised to a small and constant value and 500 extra episodes are run letting Q evolve accordingly. At each episode the corresponding policy is stored. As reported in table 1, the value of $\alpha$ and $\epsilon$ for this “policy exploration” phase are set differently according to the learning architecture considered with the purpose of exploring (sub) optimal policies without disrupting significantly the plateau that was reached by $G$. The value of $\alpha$ and $\epsilon$ that is needed to remain in the plateau reflects how robust the optimal policy is to perturbations, and appears correlated to the degree of stochasticity previously discussed. Centralized architectures appear to need stronger exploration of the plateau to “jump” between sub-optimal policies. We find a fairly stable plateau region in the exploration phase of the fully centralized 1 CC architecture even for large values of $\alpha$ and $\epsilon$. This is due to the high degeneracy of this architecture, where many different policies perform similarly, as discussed later (see fig. 5). Finally, we find that the hive update introduces some degree of stabilization of the learning process since we observed the need of less integration steps (as expected) and a smaller sensitivity to scheduling and exploration parameters. The source code for all different architectures is available at [21].
@@ III Results
Using the learning protocol described above, we store 500 optimized policies for each architecture and quantify performance as the average speed of the center of mass of the crawler for the best 250 policies. We obtained velocities by playing each policy over a trajectory of 20000 steps. In computing the average we accounted for the multiplicity of each policy, so that policies which are found several times will be more represented as well as quasi-degenerate policies realizing similar crawling velocities.
We find that performance for all architectures depends on the number of suckers $N_{s}$ (fig. 3), qualitatively consistent with the infinite continuous crawler case. Performance peaks for $N_{s}\approx 12$ and then exhibits a long-range decay beyond the peak (as discussed in Methods, the number of suckers is identical to the spatial periodicity of the CPG wave, $N$). The position of the peak depends on the physics of the crawler; in the infinite crawler it depends on the wave frequency $\omega$, see eq. 5. To investigate the limits and properties of multiagency and centralization in cooperative/coordinated actions we will focus largely on the peak. We tuned parameters so that the peak occurs at $N_{s}=12$, which is computationally treatable for all learning architectures from the most distributed to the most centralized. Whether and how natural crawlers orchestrate the central pattern generation and the suckers to operate near the peak is a fascinating question for further research.
We find that the tail and head of our finite crawler play a key role, thus distinguishing our problem from the infinite crawler for which an asymptotic solution is available. However, the analytical optimal policy is a useful heuristic for the internal suckers as discussed below. We thus construct a reference heuristic policy prescribing that (i) internal suckers follow the analytical optimal policy eq. 3 with $\beta=3/2\pi$ from [15] (see [Supplementary Video 1]), and (ii) tail and head anchor in response to negative forces, i.e. the head adheres when it is dragged backward by the elongated spring to its left and the tail adheres when it is pushed backward by the compressed spring to its right. It is important to observe that to implement the heuristic strategy, the simulator prescribes adhesion according to a phase shift, thus measuring time and space. In contrast, because we use binary compressed/elongation spring states, our agents (the suckers, or the CCs) have no access to the CPG’s wave phase. Thus, our simple crawlers cannot implement the heuristic policy, due to their limited states and controls. Intriguingly, our poor state representation can be compensated by a more centralized control which recovers a smooth traveling wave by controlling multiple suckers at a time, as we will see in the following.
@@ III.1 Distributed control
Let us first focus on distributed forms of control, where each sucker is an agent (fig. 3, left). When suckers play random actions, the center of mass does not move on average, as expected (random policy, green dashed line). If tail and head adhere in response to negative forces, even if the internal suckers perform a random action, the center of mass does move, showing the importance of the crawler’s tail and head (semi-random policy, orange dashed line). If tail and head adhere in response to negative forces and the internal suckers adhere according to the analytical results described above, performance further increases (heuristic benchmark, gray dashed line). Although the optimal policy of the internal suckers is obtained in the infinite continuous crawler limit, the clear improvement of the policy relative to the semi-random policy suggests that away from the boundary, the infinite crawler may be a useful approximation of our finite discrete crawler. When we let agents act according to a policy that has been learned through trial and error, performance is intermediate between the semi random policy and the heuristic benchmark (blue circles and red triangles). We find that standard learning is harder than hive learning, but it achieves better performance and nearly matches the heuristic benchmark described above (e.g. for 15 suckers). As a reminder, in hive learning all suckers learn the same policy, whereas in standard learning each sucker learns its own individual policy.
@@
Almost invariably trainings with hive architectures learn the same policy $\pi_{H}$, regardless of $N_{s}$. According to $\pi_{H}$, internal suckers adhere in response to the $\leftarrow$|$\leftarrow$ | ← state while the tail and head –which are connected to a single spring– adhere in response to the states ×$\times$|$\leftarrow$ | ← and |$\leftarrow$|$\times$ | × respectively. The interpretation of the hive policy $\pi_{H}$ is simple: each sucker adheres in response to negative forces, i.e. when all adjacent springs push it backward. For $N_{s}\geq 15$, all our trainings converge to the same policy $\pi_{H}$, with no suboptimal policies ever learned (hence the absence of the errorbar in fig. 3). The same behavior was found when training distributed hive policies for even larger crawlers with $N_{s}>30$ (results not shown) further confirming that $\pi_{H}$ is the best consensus policy. Departures from this policy emerge only for small $N_{s}$ where the specific dynamics of few suckers might affect the global consensus. For $N_{s}<15$ trainings converge to a small number of policies; the best learned policy is $\pi_{H}$ for $N_{s}=5$ and $12$, and it is a slightly different policy $\hat{\pi}_{H}$ only for $N_{s}=8$ and $10$. $\pi_{H}$ and $\hat{\pi}_{H}$ are almost identical, except internal suckers also adhere in response to the $\rightarrow$|$\rightarrow$ | → state, and the tail never adheres. For small number of suckers, $\hat{\pi}_{H}$ is marginally faster than $\pi_{H}$ (e.g.  for 10 suckers speed is $\approx 0.017$ vs $\approx 0.015$). Thus collective hive learning is highly transferrable, in that the hive policy can be applied to crawlers with any number of suckers with no need for retraining, consistent with previous results obtained with genetic algorithms [14]. Adopting the hive policy results in a clear traveling wave of adhesion, with only one sucker adhering at a time (see animation of the hive policy in [Supplementary Video 2]). Note that, as already remarked above, the individual suckers do not have access to their phase shift, and as a result, this architecture cannot learn a policy that implements our heuristic benchmark. Indeed, this heuristics requires a finer proprioceptive signal, e.g. inferring the phase from the precise intensity of the restoring forces, rather than its sign.
@@
Intriguingly, the standard learning, where each sucker acts according to its own policy, realizes a more complex dynamics with multiple suckers adhering simultaneously and no visible traveling wave of adhesion. In doing so it achieves a more complex group behavior and better performance. Note that, although standard learning achieves better performances, we find that training is more computationally expensive and in our hands was limited to $N_{s}=15$: beyond this limit the training did not converge. This is likely due to the intrinsic level of stochasticity: all agents learn simultaneously, but rewards are coupled, as the motion of the tentacle depends on the full dynamical state of the system. Thus good actions may be hard to learn as they may not be individually good enough to yield a positive reward. Moreover, in all cases where standard learning was successful, we noticed an extreme sensitivity to the scheduling parameters, with convergence only using about three times the number of steps per episode relative to the hive solution. In contrast, the distributed hive policies were easily learned.
@@
Figure 4: Robustness analysis of the best policy for 12 suckers crawlers with various learning architectures. a) Performance (velocity of the crawler) vs number of failing suckers. b) Performance of the optimal policy with 1 failing sucker relative to optimal performance with no failing suckers. c) Frequency each sucker is the dominant contributor to the drop in performance in a trajectory, showing that the head is the most important sucker. d) Performance drop due to failure of the head. e) Statistical behavior of a 12 sucker crawler for different distributed and centralized architectures. Top row: for each sucker (x axis), and each compression state of its adjacent springs (y axis), the color represents the frequency that state was visited while performing the optimal policy. Bottom row: Fraction of time a sucker adheres, colorcoded from purple (always adhere) to yellow (never adhere) for each of the 12 suckers (x axis) and each of the compression states of their adjacent springs (y axis). The blank entries for head and tail represent that these suckers only have access to one spring, hence two compression states (rather than four).

### v3

@@ II.4.2 Centralized control
We introduce centralization by defining an agent as an assembly of several suckers and springs which we call Control Center (CC), as illustrated by panel c of fig. 1. The number of CCs controlling the crawler depends on how many suckers are included in each CC and can range from several CCs down to a single CC, which corresponds to the highest degree of centralization. Centralized architectures control several suckers and springs simultaneously, hence have access to a richer representation of the dynamical state of the crawler, and can realize many more policies than those realizable in the distributed setting. The state of the CC is given by the compression/elongation state of each spring located between the first and last sucker within the CC; while the action of the CC consists of specifying the binary adhesion/no-adhesion instruction for each of the suckers within the CC. Since the CC has access to all combinations of compression/elongation and adhesion/no-adhesion states and actions, the state-action space scales exponentially as $|\mathcal{S}|=2^{n_{s}-1}$ and $|\mathcal{A}|=2^{n_{s}}$, with $n_{s}$ the number of suckers controlled by the given CC (for $n_{s}$ suckers under a CC there are $n_{s}-1$ connected springs). Note that to facilitate comparisons and have the same dimensionality of the Q-matrix for each CC we chose to ignore the spring between two adjacent CCs. This is the only consistent choice for hive updates, as if the spring at the interface of two control centres was included in the state, the information would be interpreted differently by the two adjacent control centres. To make a fair comparison, we keep the same choice even when each control centre learns its own policy (non hive). We do not expect considerable differences because the control centres already perform as well as the single agent, to within the error bar.
II.4.3 The hive assumption
@@
Whenever multiple agents control crawling, a simplifying assumption may be adopted imposing that all agents learn the same policy. This is the so called hive assumption or Population-Based Training [52, 43, 60], where a population of agents is used to find a shared solution. In all our multiagent learning schemes –including distributed and centralized with multiple CCs– we compare standard learning to hive learning, as illustrated in fig. 1. In practice, each agent writes and reads the same Q-matrix. The distributed hive Q-matrix becomes a 8X2 tensor, where in addition to the 4 states associated to the internal sucker there are 2 extra elements of the matrix for the tail, (null,elongated) and (null, compressed), and for the head, (elongated,null) and (compressed,null) [^4] Note that in this case the problem is effectively hive only for the internal suckers, whilst tail and head remain genuinely independent (since their states are unique) as in standard multiagent RL.. For the hive CCs there is no need to modify the Q-matrix between hive and standard learning, since all CCs have Q-matrices that are identical in shape, as discussed above. Clearly, the hive assumption speeds up training as it allows to parallelize the learning process. However, it also affects significantly the learned policies, as all agents are forced to learn the same policy, despite being placed at different locations relative to the propagating wave. In other words, the hive update imposes a consensus mechanism and all agent must “agree” on the same policy. In fact in the hive policy we find that different agents will learn distinct policies, and the hive update reduces the number of policies that are learned, as discussed in the Results section.
II.5 Learning protocol
Figure 2: Illustration of the training protocol on a 12 suckers crawler for various learning architectures. (a) Average return $G=1/|\mathcal{S}|\sum_{s}\max_{a}\left(Q(s,a)\right)$ as a function of learning episode for distributed learning architectures with the same Q matrix for all agents (hive, top) or one Q matrix per agent (standard, bottom). Top left: $G$ initially increases and then slows down; the process is repeated increasing the number of steps per episode (color lines) until learning reaches a flat plateau (black). Right: at the plateau, identified automatically (red dashed line), $\epsilon$ and $\alpha$ are increased again within an exploration phase that stores a set of 250 policies (each color represents a different policy). (b) Same as (a), for centralized architectures with 2 control centers hive (top) and standard (bottom) and for a single control center (bottom, right). See table 1 for details on the values of the parameters.
@@ II.5 Learning protocol
RL theory and algorithms are extensively covered in specialized books and articles (e.g. [49]). As briefly summarized before, in tabular Q-learning during training we iteratively update the Q-matrix using eq. 7. To keep a certain degree of exploration during training we adopt the standard procedure of following an $\epsilon$-greedy policy, that is, with a finite probability $\epsilon$ a random action is chosen instead of the one prescribed by the current policy (i.e. the action maximizing the current estimate of the value). Duriring training, both the exploration $\epsilon$ and learning rate $\alpha$ in eq. 7 are scheduled to decrease linearly, to ensure convergence to a final estimate of Q. Note that convergence to the optimal Q matrix is only guaranteed asymptotically, hence different trainings will lead to different sub-optimal policies, which we explore systematically in the following. In this work the scheduling process starts at a value of $\epsilon_{\max}=0.9$ and $\alpha_{\max}=0.1$ and is scheduled down to $\epsilon_{\min}=0.001$ and $\alpha_{\min}=0.001$. Because here we seek fast unidirectional motion, the task is continuing and there is no termination. With the term episode we intend a series of updates characterized by the same value of the parameters $\alpha$ and $\epsilon$. Note that this definition is different from the usual definition used in episodic tasks, where the end of the episode corresponds to the agent reaching a specific terminal state.
Note that although the equations of motion are deterministic, our learning problem is stochastic. There are two sources of stochasticity: i. the coarse grained state space and ii. the definition of agents. Both i. and ii. imply that the reward signal is a stochastic variable. Indeed, the speed of the crawler depends on the full dynamical state of the system, defined by the (continuous) speed, position and adhesion of all suckers and on the relaxed length of all springs. However, as per i., our crawler does not observe the full dynamical state of the system, but rather its coarse projection onto discrete binary states of compression vs elongation of all the springs. Thus given the actions of all suckers, the center of mass is not fully determined yielding a stochastic reward $R(t)$. Moreover, in all multiagent architectures each agent is unaware of other agents’ actions, as per ii., hence once again the actions of any individual agent do not individually determine the speed of the center of mass (i.e. the reward). This latter source of noise in the reward is bypassed by the fully centralized architecture with a single control center (1CC).
@@
To monitor progress during learning we consider the average value $G$, with $G=1/|\mathcal{S}|\sum_{s}\max_{a}\left(Q(s,a)\right)$. Clearly, when $G$ increases, the $Q$ matrix is being updated and learning is in progress, whereas when $G$ reaches a plateau, the learning process has come to an end. As illustrated by fig. 2, we set an automatic protocol consisting in repeating the learning procedure until a plateau in $G$ is detected: when the plateau is not encountered after 1500 episodes, we repeat the process changing the scheduling by increasing the number of steps per episode. When several Q-matrices are learned simultaneously (for instance, in standard distributed control) the convergence is monitored separately for each agent. Due to the stochastic nature of the problem, we noticed that repeating the learning procedure yields different policies. Therefore, we design a protocol to account for the stochasticity of this learning process. To efficiently store several sub-optimal policies, without repeating the entire training procedure, we introduced the following strategy: when convergence is detected, the learning parameters are raised to a small and constant value and 500 extra episodes are run letting Q evolve accordingly. At each episode the corresponding policy is stored. As reported in table 1, the value of $\alpha$ and $\epsilon$ for this “policy exploration” phase are set differently according to the learning architecture considered with the purpose of exploring (sub) optimal policies without disrupting significantly the plateau that was reached by $G$. The value of $\alpha$ and $\epsilon$ that is needed to remain in the plateau reflects how robust the optimal policy is to perturbations, and appears correlated to the degree of stochasticity previously discussed. Centralized architectures appear to need stronger exploration of the plateau to learn distinct sub-optimal policies. We find a fairly stable plateau region in the exploration phase of the fully centralized 1 CC architecture even for large values of $\alpha$ and $\epsilon$. This is due to the high degeneracy of this architecture, where many different policies perform similarly, as discussed later (see fig. 5). Finally, we find that the hive update introduces some degree of stabilization of the learning process since we observed the need of less integration steps (as expected) and a smaller sensitivity to scheduling and exploration parameters. The source code for all different architectures is available at [15]. Next we discuss results obtained by repeating the protocol described above for several centralizes and distributed architectures and a number of suckers that ranges from $N_{s}=5$ to $N_{s}=30$. We reserve a particularly in-depth analysis to the case of $N_{s}=12$.
@@ III Results
Using the learning protocol described above, we store 500 optimized policies for each architecture and quantify performance as the average speed of the center of mass of the crawler for the best 250 policies. We obtained velocities by playing each policy over a trajectory of 20000 steps, or about 300 full cycles of the CPG travelling wave across the full crawler. In computing the average we accounted for the multiplicity of each policy, so that policies which are found several times will be more represented as well as quasi-degenerate policies realizing similar crawling velocities.
We find that performance for all architectures depends on the number of suckers $N_{s}$ (fig. 3), qualitatively consistent with the infinite continuous crawler case previously discussed. With our choice of parameter, performance peaks for $N_{s}\approx 12$ and then exhibits a long-range decay beyond the peak (as discussed in Methods, the number of suckers is identical to the spatial periodicity of the CPG wave, $N$). To investigate the limits and properties of multiagency and centralization in cooperative/coordinated actions we will focus largely on the peak. The value of the parameters we chose correspond to the peak occurring near $N_{s}=12$, which is computationally treatable for all learning architectures from the most distributed to the most centralized. Whether and how natural crawlers orchestrate the central pattern generation and the suckers to operate near the peak is a fascinating question for further research.
We find that the tail and head of our finite crawler play a key role, thus distinguishing our problem from the infinite crawler for which an asymptotic solution is available. However, the analytical optimal policy is a useful heuristic for the internal suckers as discussed below. We thus construct a reference heuristic policy prescribing that (i) internal suckers follow the analytical optimal policy eq. 5 from [51] (see [Supplementary Video 1]), and (ii) tail and head anchor in response to negative forces, i.e. the head adheres when it is dragged backward by the elongated spring to its left and the tail adheres when it is pushed backward by the compressed spring to its right. It is important to observe that to implement the heuristic strategy, the simulator prescribes adhesion according to a phase shift, thus measuring time and space. In contrast, because we use binary compressed/elongation spring states, our agents (the suckers, or the CCs) have no access to the CPG’s wave phase. Thus, our simple crawlers cannot implement the heuristic policy, due to their limited states and controls. Intriguingly, our poor state representation can be compensated by a more centralized control which recovers a smooth traveling wave by controlling multiple suckers at a time, as we will see in the following.
@@ III.1 Distributed control
Let us first focus on distributed forms of control, where each sucker is an agent (fig. 3, left). When suckers play random actions, the center of mass does not move on average, as expected (random policy, green dashed line). If tail and head adhere in response to negative forces, even if the internal suckers perform a random action, the center of mass does move, showing the importance of the crawler’s tail and head (semi-random policy, orange dashed line). If tail and head adhere in response to negative forces and the internal suckers adhere according to the analytical results described above, performance further increases (heuristics, gray dashed line). Although the analitically optimal policy of the internal suckers is obtained in the infinite continuous crawler limit, the clear improvement of the policy relative to the semi-random policy suggests that away from the boundary, the infinite crawler may be a useful approximation of our finite discrete crawler. This is consistent with the fact that our crawler translocates with a speed that varies with the number of suckers, qualitatively similar to the infinite crawler in Ref. [51]. We report this heuristics as a qualitative reference to guide the eye. Note that experiments on leech [47] support the role of head and tail postulated here. When we let agents act according to a policy that has been learned through trial and error, performance is intermediate between the semi random policy and the heuristics (fig. 3a, blue circles and red triangles). We find that standard learning is harder than hive learning, but it achieves better performance and nearly matches the heuristics described above (e.g. for 15 suckers). As a reminder, in hive learning all suckers learn the same policy, whereas in standard learning each sucker learns its own individual policy.
@@
Almost invariably trainings with hive architectures learn the same policy $\pi_{H}$, regardless of $N_{s}$. According to $\pi_{H}$, internal suckers adhere in response to the $\leftarrow|\leftarrow$ state while the tail and head –which are connected to a single spring– adhere in response to the states $\times|\leftarrow$ and $\leftarrow|\times$ respectively. The interpretation of the hive policy $\pi_{H}$ is simple: each sucker adheres in response to negative forces, i.e. when all adjacent springs push it backward. For $N_{s}\geq 15$, all our trainings converge to the same policy $\pi_{H}$, with no suboptimal policies ever learned (hence the absence of the errorbar in fig. 3a). The same behavior was found when training distributed hive policies for even larger crawlers with $N_{s}>30$ (results not shown) further confirming that $\pi_{H}$ is the best consensus policy. Departures from this policy emerge only for small $N_{s}$ where the specific dynamics of few suckers might affect the global consensus. For $N_{s}<15$ trainings converge to a small number of policies; the best learned policy is $\pi_{H}$ for $N_{s}=5$ and $12$, and it is a slightly different policy $\hat{\pi}_{H}$ only for $N_{s}=8$ and $10$. $\pi_{H}$ and $\hat{\pi}_{H}$ are almost identical, except internal suckers also adhere in response to the $\rightarrow|\rightarrow$ state, and the tail never adheres. For small number of suckers, $\hat{\pi}_{H}$ is marginally faster than $\pi_{H}$ (e.g.  for 10 suckers speed is $\approx 0.017$ vs $\approx 0.015$). Thus collective hive learning is highly transferrable, in that the hive policy can be applied to crawlers with any number of suckers with no need for retraining, consistent with previous results obtained with genetic algorithms [18]. Adopting the hive policy results in a clear traveling wave of adhesion, with only one sucker adhering at a time (see animation of the hive policy in [Supplementary Video 2]). Note that, as already remarked above, the individual suckers do not have access to their phase shift, and as a result, this architecture cannot learn a policy that implements our heuristic benchmark. Indeed, this heuristics requires a finer proprioceptive signal, e.g. inferring the phase from the precise intensity of the restoring forces, rather than its sign.
@@
Intriguingly, the standard learning, where each sucker acts according to its own policy, realizes a more complex dynamics with multiple suckers adhering simultaneously and no visible traveling wave of adhesion. In doing so it achieves a more complex group behavior and better performance. Note that, although standard learning achieves better performances, we find that training is more computationally expensive and in our hands was limited to $N_{s}=15$: beyond this limit the training did not converge. This is likely due to the intrinsic level of stochasticity: all agents learn simultaneously, but rewards are coupled, as the motion of the crawler depends on the full dynamical state of the system. Thus good actions may be hard to learn as they may not be individually good enough to yield a positive reward. Moreover, in all cases where standard learning was successful, we noticed an extreme sensitivity to the scheduling parameters, with convergence only using about three times the number of steps per episode relative to the hive solution. In contrast, the distributed hive policies were easily learned.
@@
Figure 4: Robustness analysis of the best policy for 12 suckers crawlers with various learning architectures. a) Performance (velocity of the crawler) vs number of failing suckers. b) Performance of the optimal policy with 1 failing sucker relative to optimal performance with no failing suckers. c) Frequency each sucker is the dominant contributor, measured by the largest drop in performance when the sucker fails: the head is the most important sucker. d) Performance drop due to failure of the head. e) Statistical behavior of a 12 sucker crawler for different distributed and centralized architectures. Top row: for each sucker (x axis), and each compression state of its adjacent springs (y axis), the color represents the frequency that state was visited while performing the optimal policy. Bottom row: Fraction of time a sucker adheres, colorcoded from purple (always adhere) to yellow (never adhere) for each of the 12 suckers (x axis) and each of the compression states of their adjacent springs (y axis). The blank entries for head and tail represent that these suckers only have access to one spring, hence two compression states (rather than four).